OpenStudy (anonymous):

A long straight wire carries current of 6A. The magnetic flux density at a point 6cm from the conductor is .. ?

OpenStudy (shane_b):

Just use the formula you should have been given in class or in the book:\[\large B=\frac{u_0I}{2\pi d}=\frac{(1.257x10^{-6}N/A^2)(6A)}{2\pi(0.06m)}\]